Solve the system of equations
-x + 6y = -6
3x − 29y = 28
x =
y =

Answer by stanbon(75887)

Solve the system of equations
-x + 6y = -6
3x − 29y = 28
----
Multiply thru the 1st eq. by 3:
-3x + 18y = -18
3x - 29y = 28
----------
Add and solve for "y":
-11y = 10
y = -10/11
--------------
solve for "x":
-x + 6y = -6
-x + 6(10/11) = -(66/11)
-x = -126/11
x = 126/11
=====================
cheers,
Stan H.
==============
x =
y =